A small, 200 g cart is moving at 1.80 m/s on a frictionless track when it collides with a larger, 5.00 kg cart at rest. After the collision, the small cart recoils at 0.850 m/s .
Answer #1

Conservation of momentum

m1u1+m2u2=m1v1+m2v2
=>0.2 x 1.80 +m2 x 0 = 0.2 x (-0.850) + 5 x v2
=>0.36 = -0.17 + 5v2
=>5v2 = 0.36 + 0.17
=>v2 = 0.53/5
=>v2 = 0.106 m/s